The resultant of two forces acting at an angle of 120° is 10 kg wt and is perpendicular to one of the forces.what is the magnitude of that force?

Grade:11

3 Answers

Arun
4 years ago
Dear student
 
tan 30 = 1/ sqrt(3) = x/10
 
hence x = 10/sqrt(3)
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Saurabh Koranglekar
4 years ago
Dear student

Break the forces in components

F1 be on +ve axis and f2 be in 2nd quadrant, fr be perpendicular to f1

F1= f2 cos 120 = -f2/2

F2 sin120 = fr

Now f2 = 20/sqrt(3)

F1= 10/sqrt(3)

Regards
Khimraj
4 years ago
Answer is 5
Bcs if theta = 120 degree then 
Resultant=A+B 
Where 
A→=B→=a
a+a=10
a=5
..............................................